# Large Language Model Agent for Modular Task Execution in Drug Discovery

### v2

@@ 3.3 Seed Molecule Generation
The initial library of molecules serves as a critical foundation for exploring chemical space and identifying candidates for further optimization. To construct this seed molecule pool, AgentD leverages two complementary generation strategies. REINVENT enables de novo molecule generation, allowing for broad and unbiased exploration of chemical space without the need for an input structure. In contrast, Mol2Mol performs conditional generation, producing analogs that are structurally similar to a specified input molecule. In our workflow, the existing drug identified during the data extraction phase, such as venetoclax, is used as input for Mol2Mol, enabling the agent to focus on chemically relevant and biologically meaningful regions of the search space.
@@
Figure 5: t-SNE visualization of chemical space, based on Mordred-derived molecular descriptors 55, including seed molecules generated by AgentD (Mol2Mol, REINVENT), refined molecules after property-aware optimization, 5,000 randomly sampled compounds from ChEMBL, and the reference drug venetoclax.
@@ 3.4 Property-Aware Molecular Refinement
To enhance the quality of the seed molecule pool, AgentD performs property-driven refinement of SMILES structures by identifying and addressing limitations that compromise their drug-likeness. Using the Deep-PK model API 47, the agent predicts 67 ADMET properties including absorption (e.g., Caco-2 $\log P_{\text{app}}$), distribution (e.g., blood–brain barrier permeability), metabolism (e.g., CYP1A2 inhibition), excretion (e.g., drug half-life), and toxicity (e.g., biodegradability, general toxicity, and logD), as well as 7 general physicochemical properties such as lipophilicity and solubility. These properties are critical for evaluating pharmacokinetic profiles and safety. A complete list of predicted ADMET properties is provided in Table S2. In parallel, AgentD uses the BAPULM model 48 to estimate binding affinity (pKd) from the ligand SMILES and the protein FASTA sequence. Details of both models are described in the Methods section.
Based on these predictions, AgentD identifies unfavorable molecular properties relevant to drug development, such as low permeability or high toxicity, and proposes targeted structural edits to improve them. A comprehensive list of risk-associated properties is provided in Table S3. We demonstrate this refinement process over two iterations, beginning with an initial set of 100 molecules. These iterations produced 99 and 95 additional valid SMILES, respectively, which were added to the candidate pool. To illustrate how the refinement works in practice, Figure 6a shows a molecule initially flagged for poor permeability due to a low predicted $\log P_{\text{app}}$ value. In response, AgentD suggested replacing a hydroxyl group with a methyl group in the first round, followed by substituting a sulfonamide with an amine in the second - both changes contributed to improved predicted permeability.
@@
Figure 6: Representative examples of property-aware molecular refinement. Each case shows structural modifications aimed at improving a specific target property (indicated in parentheses), as proposed by the agent’s reasoning output.
@@
Quantitatively, in the first refinement round, 57 out of the initial 100 molecules were identified as having low permeability. Among these, approximately 44% showed improved $\log P_{\text{app}}$ values after refinement, 26% declined, and 30% remained unchanged. In the second round, 52% improved, 17% declined, and 31% were unchanged. Toxicity improvements were relatively more limited: 24% and 20% of high-toxicity molecules showed improvement in the first and second rounds, respectively, with most remaining toxic despite modification.
@@
Despite these local failures, the overall impact of refinement is positive across the molecule pool. This is supported by increased QED (Quantitative Estimate of Drug-likeness) scores, which reflect overall drug-likeness based on empirical physicochemical properties 56. As shown in Figure 7c, the distribution of high-QED molecules expanded with each refinement iteration: the number of molecules with QED >> 0.6 increased from 34 in the original set to 49 after the first update, and to 55 after the second.
@@
As shown in Figure 7d, LLM-based refinement increased the proportion of molecules meeting empirical drug-likeness and lead-likeness criteria, including Lipinski’s Rule of Five, Veber’s rule, Ghose’s filter, and Oprea’s lead-likeness filter 57, 58, 59, 60. These rules assess different aspects of molecular suitability: Lipinski’s and Veber’s rules emphasize oral bioavailability and permeability, Ghose’s filter captures physicochemical ranges observed in marketed drugs, and Oprea’s filter identifies compounds suitable as lead-like starting points for optimization. After refinement, more molecules satisfied Veber’s rule (78 to 86) and Oprea’s filter (60 to 67). The most notable improvement was observed for Ghose’s filter, with the number of passing molecules increasing from 32 to 55 ( 72% increase). These results indicate that refinement helped expand the chemical space toward drug-like and lead-like regions, particularly in terms of Ghose’s statistical descriptors. Compliance with Lipinski’s rule remained largely unchanged; this outcome is further discussed in the Discussion section.
Figure 7: Property changes across iterations of property-aware molecular refinement. a and b show the proportion of molecules with improved, declined, or unchanged permeability (logPapp) and toxicity-related properties. The number of molecules targeted in each refinement round is indicated in the legend. c Kernel density plot showing the distribution shift in QED scores across refinement iterations. d Histogram illustrating the number of molecules satisfying empirical drug-likeness rules; detailed rule definitions are provided in the Methods section.
@@
After refining the seed molecule pool through property-aware modifications, we applied empirical filtering criteria to select candidates for 3D structural evaluation. This filtering scheme is intended as a demonstration rather than a definitive selection strategy. Its purpose is not to identify a final drug candidate, but to illustrate the agent’s capability to perform key tasks in the drug discovery pipeline. In our workflow, molecules are shortlisted if they (i) satisfy Oprea’s lead-likeness filter, (ii) pass at least two of the three drug-likeness rules (Lipinski’s Rule of Five, Veber’s rule, and Ghose’s filter), (iii) achieve a QED score above 0.55, and (iv) have a predicted pKd value greater than 6.0.
@@
The molecule shown in Figure 8 satisfies the lead-likeness requirement, passes all three drug-likeness rules, has a QED score of 0.68, and achieves a predicted pKd of 6.18, making it a representative candidate for further analysis. It is important to note that these empirical rules may not always align, and molecules rarely satisfy all of them simultaneously. In fact, only one molecule in our set passed all three rules, which we selected as an illustrative example for 3D visualization. Additional examples are provided in Supplementary Figure S4.
Once candidates are selected, AgentD uses Boltz-2 to generate 3D protein–ligand complex structures using the ligand’s SMILES and the target protein’s FASTA sequence. These structures provide a foundation for more rigorous downstream evaluations, such as molecular docking, MD simulations, and free energy perturbation analyses, to assess binding stability, conformational flexibility, and interaction specificity under biologically relevant conditions. In addition to structural outputs, Boltz-2 returns estimated binding metrics, including logIC50 values and inhibitor probabilities. In the example shown, the predicted IC50 (0.54 µM) and inhibitor probability indicate good sub-micromolar binding affinity, although further molecular optimization or experimental validation may still be necessary.
@@
Figure 8: Protein–ligand structure generation and evaluation. a, A representative 3D complex predicted by Boltz-2, with estimated IC50 and inhibitor probability. b, Radar plot showing empirical rule-based evaluation. Green dotted lines indicate the thresholds for each rule, while the purple solid line represents the properties of the selected molecule. See Methods and Table S4 for criteria.
@@ 4 Discussion
The agent reliably retrieves protein and ligand data from both structured databases and web sources. While our current implementation focuses on targeted extraction, future extensions could incorporate large-scale literature mining and relevance scoring to support broader knowledge synthesis. For domain-specific question answering, AgentD’s RAG responses consistently outperform standard LLM outputs in contextual accuracy and mechanistic detail. However, since RAG performance strongly depends on the quality of retrieved literature, extending access beyond open-access sources could further improve its effectiveness.
@@
The property-aware molecular refinement module shows both promise and limitations. AgentD successfully improves target properties such as permeability or toxicity using only its built-in reasoning and domain knowledge. However, molecular optimization remains intrinsically complex due to interdependencies among properties. For example, in Figure 6a and b, predicted apparent permeability (logPapp) is influenced by lipophilicity (logP), consistent with Overton’s rule 61, whereas aqueous solubility (logS) does not follow a consistent trend and may either increase or decrease. These observations highlight the complex, non-linear interplay among lipophilicity, solubility, and permeability. Similarly, the first-round refinement in Figure 6d mitigates predicted liver toxicity but worsens permeability. These trade-offs highlight the fundamental challenge of multi-property optimization in drug design. These findings highlight the importance of incorporating multi-objective optimization in future iterations to better balance competing pharmacological goals.
Despite these complexities, AgentD achieves a net improvement in overall drug-likeness across the molecule pool, as shown in Figure 7c and d. Notably, the most significant improvement is observed for the Ghose filter. Because the Ghose filter reflects physicochemical property ranges derived from known drugs, this trend suggests that the LLM may leverage its training knowledge of existing molecules when refining candidates. In contrast, performance on Lipinski’s Rule of Five slightly deteriorates. Since this rule emphasizes oral bioavailability, and oral absorption was rarely designated as a primary weakness in the molecular refinement tasks (Table S3), this limited improvement is not unexpected.

### v3

@@ 3.3 Seed Molecule Generation
The initial library of molecules serves as a critical foundation for exploring chemical space and identifying candidates for further optimization. To construct this seed molecule pool, AgentD leverages two complementary generation strategies. REINVENT enables de novo molecule generation, allowing for broad and unbiased exploration of chemical space without the need for an input structure. Mol2Mol performs conditional generation, producing analogs that are structurally similar to a specified input molecule. In our workflow, the existing drug identified during the data extraction phase, such as venetoclax, is used as input for Mol2Mol, enabling the agent to focus on chemically relevant and biologically meaningful regions of the search space.
Figure 4: t-SNE visualization of chemical space, based on Mordred-derived molecular descriptors mordred, including seed molecules generated by AgentD (Mol2Mol, REINVENT), refined molecules after property-aware optimization, 5,000 randomly sampled compounds from ChEMBL, and the reference drug venetoclax.
@@ 3.4 Property-Aware Molecular Refinement
To enhance the quality of the seed molecule pool, AgentD performs property-driven refinement of SMILES structures by identifying and addressing limitations that compromise their drug-likeness. Using the Deep-PK model API deep_pk, the agent predicts 67 ADMET properties, as well as 7 general physicochemical properties. A complete list of predicted ADMET properties is provided in Supplementary Table S2. In parallel, AgentD uses the BAPULM model bapulm to estimate binding affinity (pKd).
Based on these predictions, AgentD identifies unfavorable molecular properties relevant to drug development, such as low permeability or high toxicity, and proposes targeted structural edits to improve them. A comprehensive list of weakness-associated properties is provided in Supplementary Table S3. We demonstrate this refinement process over two iterations, beginning with an initial set of 100 molecules. These iterations produced 99 and 95 additional valid SMILES, respectively, which were added to the candidate pool. For instance, a molecule is initially flagged for poor permeability due to a low predicted $\log P_{\text{app}}$ value (Figure 5a). In response, AgentD suggested replacing a hydroxyl group with a methyl group in the first round, followed by substituting a sulfonamide with an amine in the second - both changes contributed to improved predicted permeability. Figure 5d illustrates another case of successful refinement, where the targeted weak properties, such as toxicity and permeability, were improved at each step.
@@
Figure 5: Property-aware molecular refinements by AgentD. a Iterative changes improve permeability. b An unintended modification leads to improved permeability. c An intended modification fails to improve permeability. d Iterative changes improve both toxicity and permeability.
@@
In the first refinement round, 57 out of the initial 100 molecules were identified as having low permeability. Among these, approximately 44% showed improved $\log P_{\text{app}}$ values after refinement, 26% declined, and 30% remained unchanged (Figure 6a). In the second round, 52% improved, 17% declined, and 31% were unchanged. For the molecules that showed target property improvement, the median $\log P_{\text{app}}$ increased from –5.33 to –5.03 (Figure 6b). Although the magnitude of change is modest, the upward trend is clearly visible within just two iterations. Toxicity improvements were relatively more limited: 24% and 20% of high-toxicity molecules showed improvement in the first and second rounds, respectively, with most remaining toxic despite modification (Figure 6c).
Figure 6: Property-aware molecular refinement results. a Proportion of molecules with improved, declined, or unchanged permeability (logPapp). b Boxplot of logPapp for entries showing improvement. c Molecular refinement patterns for molecules with improved permeability. d QED distribution shifts across iterations (vanilla GPT-4o indicates refinement solely through direct queries to GPT-4o). e Counts of molecules satisfying empirical drug-likeness rules (definitions in Methods). f Fraction of molecules modified for permeability; “Analysis failed” indicates unresolved substructure differences by the substructure algorithm. g Functional-group frequencies in input vs. refined molecules across iterations (input = input for refinement; refined = post-refinement outputs).
@@
Despite occasional local failures, as illustrated in Figure 5b and c, the overall impact of refinement is positive across the molecule pool. This is supported by increased QED (Quantitative Estimate of Drug-likeness) scores, which reflect overall drug-likeness based on empirical physicochemical properties qed. As shown in Figure 6d, the distribution of high-QED molecules expanded with each refinement iteration: the number of molecules with QED >> 0.6 increased from 34 in the original set to 49 after the first update, and to 55 after the second.
@@
As shown in Figure 6e, LLM-based refinement increased the proportion of molecules meeting empirical drug-likeness and lead-likeness criteria, including Lipinski’s Rule of Five, Veber’s rule, Ghose’s filter, and Oprea’s lead-likeness filter lipinski2001experimental, veber2002molecular, ghose1999knowledge, oprea2001lead. These rules assess different aspects of molecular suitability: Lipinski’s and Veber’s rules emphasize oral bioavailability and permeability, Ghose’s filter captures physicochemical ranges observed in marketed drugs, and Oprea’s filter identifies compounds suitable as lead-like starting points for optimization. After refinement, more molecules satisfied Veber’s rule (78 to 86) and Oprea’s filter (60 to 67). The most notable improvement was observed for Ghose’s filter, with the number of passing molecules increasing from 32 to 55 ($\sim$72% increase). These results indicate that refinement helped expand the chemical space toward drug-like and lead-like regions, particularly in terms of Ghose’s statistical descriptors. Compliance with Lipinski’s rule remained largely unchanged; this outcome is further discussed in the Discussion section.
AgentD tends to prioritize edits to polar or electron-withdrawing groups, such as hydroxyls, sulfonamides, and nitro groups, when attempting to improve permeability-related properties. Figure 6f shows the distribution of SMILES modifications for entries with permeability as the target weakness property. The most frequent edits involved hydroxyl groups (30 cases, 15.5%), followed by sulfonamide substitutions (29 cases, 15.0%) and nitro group changes (21 cases, 10.8%). Other commonly modified motifs included alkyl groups (8.3%) and halogens (6.2%), and together these five categories account for 55.7% of all refinements. Also, AgentD’s refinement behavior does not merely reflect functional-group prevalence. For instance, despite nitro groups being less common than alkyl, sulfonamide, or hydroxyl groups, nitro edits still represent a substantial portion of modifications (10.8%; Figure 6g).
@@
After refining the seed molecule pool through property-aware modifications, we applied empirical filtering criteria to select candidates for 3D structural evaluation. The purpose of this filtering scheme is not to identify a final drug candidate, but to illustrate the agent’s capability to perform key tasks in the drug discovery pipeline. In our workflow, molecules are shortlisted if they (i) satisfy Oprea’s lead-likeness filter, (ii) pass at least two of the three drug-likeness rules (Lipinski’s Rule of Five, Veber’s rule, and Ghose’s filter), (iii) achieve a QED score above 0.55, and (iv) have a predicted pKd value greater than 6.0.
@@
The molecule shown in Figure 7 satisfies the lead-likeness requirement, passes all three drug-likeness rules, has a QED score of 0.68, and achieves a predicted pKd of 6.18. This is simply a representative candidate for further analysis. It is important to note that these empirical rules may not always align, and molecules rarely satisfy all of them simultaneously. In fact, only one molecule in our set passed all three rules, which we selected as an illustrative example for 3D visualization. Additional examples are provided in Supplementary Figure S7.
@@
Figure 7: Protein–ligand structure generation and evaluation. a A representative 3D complex predicted by Boltz-2, with estimated IC50 and inhibitor probability. b Radar plot showing empirical rule-based evaluation. Green dotted lines indicate the thresholds for each rule, while the purple solid line represents the properties of the selected molecule. See Methods and Table S4 for criteria.
AgentD uses Boltz-2 to generate 3D protein–ligand complex structures using the ligand’s SMILES and the target protein’s FASTA sequence. These structures provide a foundation for more rigorous downstream evaluations, such as molecular docking, MD simulations, and free energy perturbation analyses, to assess binding stability, conformational flexibility, and interaction specificity under biologically relevant conditions. In addition to 3D structure, Boltz-2 returns estimated binding metrics, including logIC50 values and inhibitor probabilities. In the example shown, the predicted IC50 (0.54 µM) and inhibitor probability indicate good sub-micromolar binding affinity, making the molecule a suitable candidate for further validation using molecular dynamics simulations.
3.6 Comparison Across Language Models
AgentD can operate with a variety of LLMs, including those from OpenAI, Anthropic, and DeepSeek. In practice, the framework is optimized for GPT-4o. We tested AgentD using GPT-4o, Claude-3.7-Sonnet, and DeepSeek-Chat openai2023gpt4, anthropic2024claude, deepseekai2025, lu2024deepseekvl. DeepSeek-Chat, an open-source model, successfully executed all six tasks and demonstrated behavior broadly comparable to GPT-4o. In contrast, Claude-3.7-Sonnet exhibited failure modes in two tasks: domain-specific question answering and molecular refinement.
In the question-answering task, Claude-3.7-Sonnet encountered syntax errors when parsing PDF files. AgentD, powered by GPT-4o, achieved higher BERTScores than AgentD with DeepSeek-Chat (Table 2). Among the vanilla language models, DeepSeek-Chat outperformed GPT-4o, highlighting that the performance gain provided by AgentD is more pronounced when using GPT-4o.
In the molecular refinement task, AgentD with Claude-3.7-Sonnet occasionally became trapped in repeated attempts to generate valid updated SMILES, despite explicit instructions to avoid infinite retries (see Supplementary Figure S4). For the quantitative comparison of property-aware molecular refinement shown in Figure 8, we report results from a verified successful run of Claude-3.7-Sonnet to ensure a fair comparison.
Figure 8: Comparison of AgentD performance across different language models. a Distribution of QED values before and after two iterations of property-aware molecular refinement for AgentD powered by GPT-4o, Claude-3.7-Sonnet, and DeepSeek-Chat. b Histogram showing the number of molecules passing empirical lead- and drug-like rules. c–e Tanimoto similarity between the original and updated SMILES for the top five most frequently targeted weakness properties for AgentD powered by GPT-4o (c), Claude-3.7-Sonnet (d), and DeepSeek-Chat (e).
All three model cases improved the QED of seed molecules after two iterations of property-aware refinement (Figure 8a). Starting from an average QED of 0.47, AgentD, powered by GPT-4o, increased the mean value to 0.57, Claude-3.7-Sonnet increased it to 0.59, and DeepSeek-Chat increased it to 0.55. While all models demonstrated meaningful improvements, DeepSeek-Chat achieved a slightly smaller gain relative to the others. The seed set contained 23.0% of molecules with QED >0.6, which increased to 45.7% with GPT-4o, 44.9% with Claude-3.7-Sonnet, and 40.0% with DeepSeek-Chat after the second refinement step.
The rule-based evaluation shows a similar trend across the three LLMs, with the most evident increase observed in the number of molecules passing the Ghose filter (Figure 8b). Claude-3.7-Sonnet and DeepSeek-Chat do not exhibit the slight deterioration in Lipinski’s Rule of Five observed with GPT-4o, although both models yield slightly fewer molecules passing Veber’s rule.
Each language model exhibits some variability in the selection of “weakness” properties during molecular refinement. Figures 8c–e report the Tanimoto similarity tanimoto, tanimoto2 between the original SMILES and the updated SMILES for the top five most frequently targeted weakness properties. For AgentD powered by GPT-4o, the top five properties were logPapp, AMES mutagenicity, liver injury I toxicity, logP, and hERG blocker toxicity, with median similarities of 0.71, 0.71, 0.67, 0.81, and 0.79, respectively. Despite model-specific differences, the selected properties largely arise from a consistent subset of the 75 property endpoints, indicating broadly aligned refinement priorities. AgentD with Claude-3.7-Sonnet prioritized liver injury I toxicity, hERG blocker toxicity, AMES mutagenicity, liver injury II toxicity, and logP, with median similarities of 0.76, 0.79, 0.73, 0.79, and 0.83. For AgentD with DeepSeek-Chat, the top five properties were liver injury II toxicity, hERG blocker toxicity, liver injury I toxicity, AMES mutagenicity, and logPapp, with median similarities of 0.80, 0.80, 0.77, 0.70, and 0.77. Although all three models generally target similar toxicity- and permeability-related weaknesses, the specific ranking of properties and the degree of structural change.
@@ 4 Discussion
4.1 Property Awareness Improves Drug-Likeness
Weakness-property awareness substantially improves drug-likeness, even though drug-likeness is not explicitly used as an optimization target. Refinement via a vanilla LLM, without property-aware guidance, yields little measurable improvement (Figure 6d–e). In contrast, incorporating weakness-property information leads to pronounced gains, indicating that providing explicit context about what is deficient in a molecule enables more targeted and pharmacologically meaningful edits. This functions similarly to chain-of-thought reasoning: giving the LLM stepwise situational cues (e.g., “this is the weakness, so modify accordingly”) helps steer its structural decisions in a more directed and interpretable manner.
AgentD’s refinement pattern also does not simply mirror the prevalence of functional groups in the dataset (Figure 6f–g). Instead, the model displays distinct preferences in the types of chemical transformations it applies, suggesting that the observed patterns arise from its learned reasoning rather than from dataset-driven frequencies.
The benefit of property-aware refinement is consistent across all LLMs used as the agent’s core engine (Figure 8a–b). Whether using GPT-4o, Claude-3.7-Sonnet, or DeepSeek-Chat, the refinement strategy reliably increases drug-likeness. This indicates that the performance gains stem from the refinement paradigm itself rather than from model-specific idiosyncrasies.
4.2 Challenges in Molecular Refinement
The property-aware molecular refinement module also shows limitations. Molecular optimization remains intrinsically complex due to interdependencies among properties. For example, in Figure 5a and b, predicted apparent permeability (logPapp) is influenced by lipophilicity (logP), consistent with Overton’s rule overton, whereas aqueous solubility (logS) does not follow a consistent trend and may either increase or decrease. These observations highlight the complex, non-linear interplay among lipophilicity, solubility, and permeability. Similarly, the first-round refinement in Figure 5d mitigates predicted liver toxicity but worsens permeability. These trade-offs highlight the fundamental challenge of multi-property optimization in drug design. These findings highlight the importance of incorporating multi-objective optimization in future iterations to better balance competing pharmacological goals. Despite these complexities, AgentD achieves a net improvement in overall drug-likeness across the molecule pool, as shown in Figure 6c and d.
Molecular optimization is also target-dependent, as multiple interacting physicochemical properties influence outcomes. Across the three therapeutic targets, BCL-2, JAK-2, and thrombin, drug-likeness improves overall, as reflected by QED distributions and rule-based filters, though the improvement patterns differ. For BCL-2 and thrombin, the largest gains occur under the Ghose filter, while Lipinski’s Rule of Five shows only minor improvement, likely because oral bioavailability was rarely selected as a weakness (Supplementary Table S3). By contrast, JAK-2 shows stronger gains under Lipinski’s criteria but minimal change under the Ghose filter. These differences highlight the intrinsic complexity of molecular refinement, where multiple interdependent properties are convoluted.
4.3 Effective Retrieval and Generation
The agent reliably retrieves protein and ligand data from both structured databases and web sources. While our current implementation focuses on targeted extraction, future extensions could incorporate large-scale literature mining and relevance scoring to support broader knowledge synthesis.
For domain-specific question answering, AgentD’s RAG-augmented responses consistently outperform standard LLM outputs in contextual accuracy and mechanistic completeness (Table 2). Because the effectiveness of RAG depends strongly on the quality of retrieved literature, expanding retrieval beyond open-access sources may further enhance performance. Our current implementation uses a deliberately minimal retrieval pipeline, without filters on publication date, venue quality, or citation metrics, yet it still provides clear gains in factual grounding and mechanistic fidelity. These findings demonstrate that even a raw, loosely constrained RAG setup can significantly improve scientific question answering.
@@ 4.3 Effective Retrieval and Generation
4.4 Limitations and Future Directions
Despite its modular design and encouraging end-to-end performance, AgentD has several limitations. The workflow relies on external predictive tools such as Deep-PK, BAPULM, and Boltz-2, each of which carries its own error profile and uncertainties. Inaccuracies in these upstream modules can propagate downstream, affecting both refinement decisions and compound prioritization. While a fault-tolerance scheme has been implemented for Deep-PK, similar safeguards are not yet in place for other components.
Another source of variability arises from the inherently non-deterministic nature of weakness property identification during refinement. Because the agent may select different properties as optimization targets across runs, exact reproduction of a single refinement is challenging. For this reason, we analyze outcomes at the group level rather than on a per-run or per-molecule basis. Additionally, the RAG module is currently configured in a minimal form, performing unfiltered retrieval without consideration of venue quality or citation strength.
Future iterations of AgentD could address these limitations by integrating uncertainty-aware property prediction, ensemble scoring, and confidence-weighted refinement to reduce sensitivity to individual tool errors. More sophisticated retrieval strategies, such as relevance ranking or knowledge graph integration, would strengthen literature grounding, while feedback from molecular dynamics or free-energy calculations could provide physically validated guidance for structural optimization.